OpenStudy (volleyballlover55):

The mass of the Rock of Gibraltar is 1.78 ⋅ 10^12 kilograms. The mass of the Antarctic iceberg is 4.55 ⋅ 10^13 kilograms. Approximately how many more kilograms is the mass of the Antarctic iceberg than the mass of the Rock of Gibraltar? Show your work and write your answer in scientific notation. Make sure to write the steps you used to solve the problem as well! :)

OpenStudy (volleyballlover55):

@Mehek14 @mathmate @mathstudent55 @MysticFury

OpenStudy (mysticfury):

i was never good at these types of problems ......

OpenStudy (mathmate):

I'll help by giving a similar example. The mass of Mars is \(6.39\times 10^{23}\) The mass of Earth is \(5.97\times 10^{24}\) To find the difference, we have to make a common factor between the two, seeing that \(10^{24} = 10*10^{23}\) we write Difference = \(5.97*10\times 10^{23}-6.39\times 10^{23}\) =\(59.7\times 10^{23}-6.39\times 10^{23}\) =\((59.7-6.39)\times 10^{23}\) =\(53.3\times 10^{23}\) So you can proceed similarly to solve the Gibraltar problem.

OpenStudy (volleyballlover55):

OK @mathmate is this the answer? Step 1: 4.55 x 10^13 - 1.78 x 10^12 ------------------ rewrite - 45.50 x 10^12 - 1.78 x 10^12 ------------------- 43.72 x 10^12 Step 2: Then re-write it using correct scientific notation. 4.372 x 10^13.

OpenStudy (volleyballlover55):

@AnitaS. @Mehek14

OpenStudy (volleyballlover55):

@Vocaloid

Vocaloid (vocaloid):

yes

OpenStudy (volleyballlover55):

ok! :)

OpenStudy (mathmate):

@volleyballlover55 Yes, it's perfect! Sorry for being late, but luckily @vocaloid beat me to it!